THE RADIO SPECTRA OF SICCH, SICN AND SiNC

\underline{A. J. APPONI}, M. C. MCCARTHY, C. A. GOTTLIEB, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138..

Three new silicon-bearing radicals of astrophysical interest, SiCCH, SiCN and SiNC, were detected in a laboratory discharge in their X^{2}Π ground states by Fourier transform microwave and direct absorption millimeter-wave spectroscopy. Well-resolved Λ–doubling was observed in both fine structure components, and hyperfine structure was observed in the low rotational transitions of the ^{2}Π_{1/2} ladder. With the spectroscopic constants derived from the laboratory measurements, the spectra of the three radicals can be calculated to an uncertainty of less than 0.1 km s^{−1} in equivalent radial velocity over the entire range of interest to radio astronomers. SiCN with a dipole moment of 2.9 D is probably the most promising of the three for astronomical discovery.